OpenStudy (madeleine2):

In physics, if a moving object has a starting position at s 0, an initial velocity of v 0, and a constant acceleration a, then the position S at any time t > 0 is given by: S = at 2 + v 0 t + s 0. Solve for the acceleration, a, in terms of the other variables. For this assessment item, you can use ^ to show exponents and type your answer in the answer box, or you may choose to write your answer on paper and upload it.

OpenStudy (welshfella):

Sorry i misread the question its just algebra rearrange to make a the subject of the equation

OpenStudy (welshfella):

s = at^2 + v0 t + s0 subtract Vo t and s0 from both sides of the equation s - v0 t - s0 = at^2 now divide both sides by t^2 and you have the answer

OpenStudy (madeleine2):

how would i divide from both sides of the equation ? @welshfella

OpenStudy (welshfella):

divide each term by t^2 the first bit on left side would be s / t^2 then do same to other 2 terms dividing at^2 on right side by t^2 will give a , which is what you want

OpenStudy (welshfella):

s/t^2 - _____ - ______ = a can you fill in the blanks?

OpenStudy (welshfella):

just divide v0 t by t^2 and s0 by t^2

OpenStudy (madeleine2):

s/t^2-s - v0 t - s0 = at^2?

OpenStudy (welshfella):

No you haven't done the division

OpenStudy (welshfella):

I did the right side already its a

OpenStudy (madeleine2):

im so confused

OpenStudy (madeleine2):

s/t^2-s - v0 t^2 - s0t^2 = at^2

OpenStudy (welshfella):

s divided by t^2 = s / t^2 in the same way v0 t divided bt t^2 = v0 t / t^2

OpenStudy (madeleine2):

s / t^2-v0 t / t^2

OpenStudy (welshfella):

yea now do the same to s0 and at^2

OpenStudy (madeleine2):

s / t^2-v0 t / t^2 s0/t^2

OpenStudy (madeleine2):

is s / t^2-v0 t / t^2 s0/t^2 the answer?

OpenStudy (welshfella):

s - V0 t - s0 = at^2 To get a on its own on right side we divide by t^2 :- s/t^2 - V0 t / t^2 - s0 / t^2 = at^2 / t^2 Now simplify s/t^2 - v0 / t - s0/t^2 = a

OpenStudy (welshfella):

v0 t / t^2 = v0 / t beacuse t will divide into t^2 to give t on the bottom

OpenStudy (welshfella):

OK we have solved the original equation for a. a = s/t^2 - v0 / t - s0 /t^2

OpenStudy (welshfella):

- I flipped the equation over so that a is on left side.

OpenStudy (welshfella):

gotta go